Saratoga Ag Industries has 14,000,000 shares outstanding with a current stock price of $4.00. The company is going to declare a one-for-six reverse stock split. What is the market value of the company prior to the reverse stock split? What will the stock price be after the reverse split? How many shares will be outstanding after the reverse split? What is the market value after the reverse split? Numbers greater than 1,000, round to the nearest whole number and no commas. Numbers less than 1,000, round to two decimals.
Answer #1

market value = 14,000,000 * 4 = 56,000,000

stock price = 4*6 = 24

shares left = 14,000,000/6 = 2,333,333

market value remains same = 56,000,000